# Few-shot Learner Parameterization by Diffusion Time-steps

### v1

@@ 2 Related Works
FSL with Foundation Models. There are two main approaches that both leverage CLIP. First is prompt tuning, which aims to learn a prompt for each class. CoOp [47] learns a continuous prompt embedding instead of using a hand-crafted prompt. CoCoOp [46] extends CoOp by learning an image conditional prompt. ProGrad [48] aligns the prompt gradient to the general knowledge of CLIP. Recent MaPLe [14] additionally fine-tune the CLIP visual encoder. The other line aims to learn a CLIP visual feature adapter. CLIP-Adapter [7] applies a lightweight residual-style adapter, followed by the training-free approach Tip-Adapter [43]. CALIP [9] proposes a parameter-free attention to improve both zero-shot and few-shot performance. The recent CaFo [44] ensembles multiple foundation models to help with feature adaptation. However, they still suffer from the spurious correlation.
@@ 3.3 Theory
Theorem. 1) For each $\mathbf{c}\in\mathcal{C}$, there exists a smallest time-step $t(\mathbf{c})$, such that $\mathbf{c}$ is lost with at least degree $\tau$ at each $t\in\{t(\mathbf{c}),\ldots,T\}$. This also holds for each $\mathbf{e}\in\mathcal{E}$. 2) $\exists\{\beta_{i}\}_{i=1}^{T}$ such that $t(\mathbf{e})>t(\mathbf{c})$ whenever $\lVert\Phi(\mathbf{c}^{\prime},\mathbf{e})-\Phi(\mathbf{c}^{\prime},\mathbf{e}^{\prime})\rVert$ is first-order stochastic dominant over $\lVert\Phi(\mathbf{c},\mathbf{e}^{\prime})-\Phi(\mathbf{c}^{\prime},\mathbf{e}^{\prime})\rVert$ with $\mathbf{c}^{\prime}\sim\mathcal{C},\mathbf{e}^{\prime}\sim\mathcal{E}$ uniformly.
@@ 4.2 De-biasing by Time-steps
1
2
4
8
16
1
2
4
8
16
CLIP
Zero-Shot [25]
24.9 12.5
CoOp [47]
22.8 -2.1
28.4 +3.5
32.4 +7.5
37.7 +12.8
40.5 +15.6
10.8 -1.7
19.3 +6.8
19.6 +7.1
24.2 +11.7
25.8 +13.3
Co-CoOp [46]
30.1 +5.2
31.6 +6.7
33.6 +8.7
37.3 +12.4
38.2 +13.3
11.4 -1.1
13.9 +1.4
16.6 +4.1
21.8 +9.3
22.8 +10.3
MaPLe* [14]
30.1 +5.2
33.0 +8.1
33.8 +8.9
39.4 +14.5
40.7 +15.8
11.8 -0.7
14.5 +2.0
18.0 +5.5
22.6 +10.1
26.8 +14.3
OpenCLIP
Zero-Shot [25]
42.3 16.9
Linear-probe [25]
18.4 -23.9
32.5 -9.8
44.1 +1.8
55.0 +12.7
59.8 +17.5
12.4 -4.5
14.5 -2.4
17.7 +0.8
20.1 +3.2
21.6 +4.7
Tip-Adapter [43]
47.7 +5.4
51.6 +9.3
54.7 +12.4
58.4 +16.0
62.2 +19.9
22.6 +5.7
25.3 +8.4
23.6 +6.7
31.1 +14.2
33.9 +17.0
Tip-Adapter-F [43]
48.4 +6.1
53.9 +11.6
56.9 +14.7
62.0 +19.7
67.4 +25.1
21.4 +4.5
22.3 +5.4
26.8 +9.9
34.4 +17.5
40.3 +23.4
CaFo* [44]
51.7 +9.4
54.6 +12.3
58.5 +16.2
63.2 +20.9
66.7 +24.4
DM
Zero-Shot [16]
24.3 11.7
@@ 5.2 Main Results
1
2
4
8
16
1
2
4
8
16
CLIP
CoOp [47]
8.2
10.4
17.2
29.6
33.7
11.4
13.9
18.1
30.3
34.1
Co-CoOp [46]
9.7
12.1
20.1
32.7
44.5
30.1
31.6
33.6
37.3
38.2
MaPLe [14]
13.5
32.9
40.7
55.7
63.0
35.2
40.7
44.4
57.5
68.1
OC
Linear-probe [25]
11.9
13.2
37.7
53.5
60.2
16.7
29.5
51.1
61.1
69.8
Tip-Adapter [43]
28.5
36.9
46.5
59.3
66.7
36.4
47.5
59.8
71.2
80.1
Tip-Adapter-F [43]
29.3
32.0
54.4
74.2
82.6
37.4
48.6
62.2
79.1
85.4
@@
1
2
4
8
16
1
2
4
8
16
Subset
last
48.5 55.8
64.2
73.1
78.1
36.9 53.6 73.1
84.2
91.3
last + $w_{1}$
48.2
55.7
65.0
74.2
78.5
28.7
47.7
69.7
84.5
91.1
last + $w_{1}$+$w_{0}$
46.5
55.4
65.1
74.0
79.9
25.8
43.2
65.7
83.4
91.6
Weight
ELB weight
44.7
50.2
59.9
68.7
71.2
33.2
50.2
68.5
78.3
81.8
PDAE
49.2 56.7
64.1
72.2
78.1
36.2
53.9 73.4
81.2
90.8
Ours
48.5
55.8
64.2 74.2 79.9 36.9
53.6
73.1
84.5 91.6
@@
Overall Results. As shown in Table 1 and 2, our TiF learner achieves state-of-the-art performance across fine-grained and reID datasets, e.g., significantly improving existing methods by 13.7% on FGVCAircraft, 21.6% on DukeMTMC-reID and 16% on VeRi-776. On ISIC2019, we still achieve up to 7% absolute gain over the best-performing baseline, despite the challenges posed by the diverse appearances within each class, as well as the stringent macro-F1 evaluation, which is sensitive to under-performing classes. Note that prompt tuning methods generally have non-ideal performances. This holds especially on customized reID classes with no class name to provide semantic prior. In fact, this validates the difficulties to describe the specification of fine-grained classes by prompt alone.
@@ 5.3 Ablations
Figure 6: Top: Improvements in reconstruction loss by using parameterization $(\theta_{c},y,t)$ over a pre-trained SD at each $t$. Bottom: The gap in reconstruction loss of using a wrong parameterization $(\theta_{c^{\prime}},y,t)$ over using the ground-truth one $(\theta_{c},y,t)$.
@@
Parameterization Quality. We further validate our theory that connects small time-step to fine-grained attributes in Figure 6, where a small time-step corresponds to the most improvements of reconstructing fine-grained classes (after training with Eq. (6)), and also corresponds to the most discriminative one to tell the ground-truth class of each test image from the rest classes.

### v2

@@ 2 Related Works
FSL with Foundation Models. There are two main approaches that both leverage CLIP. First is prompt tuning, which aims to learn a prompt for each class. CoOp [48] learns a continuous prompt embedding instead of using a hand-crafted prompt. CoCoOp [47] extends CoOp by learning an image conditional prompt. ProGrad [49] aligns the prompt gradient to the general knowledge of CLIP. Recent MaPLe [15] additionally fine-tune the CLIP visual encoder. The other line aims to learn a CLIP visual feature adapter. CLIP-Adapter [8] applies a lightweight residual-style adapter, followed by the training-free approach Tip-Adapter [44]. CALIP [10] proposes a parameter-free attention to improve both zero-shot and few-shot performance. The recent CaFo [45] ensembles multiple foundation models to help with feature adaptation. However, they still suffer from the spurious correlation. Besides CLIP-based approaches, recent works have explored in-context learning with vision-language models [1], yet their current classification accuracy still lags behind.
@@ 3.3 Theory
Theorem. 1) For each $\mathbf{c}\in\mathcal{C}$, there exists a smallest time-step $t(\mathbf{c})$, such that $\mathbf{c}$ is lost with at least degree $\tau$ at each $t\in\{t(\mathbf{c}),\ldots,T\}$. This also holds for each $\mathbf{e}\in\mathcal{E}$. 2) $\exists T,\{\beta_{i}\}_{i=1}^{T}$ such that $t(\mathbf{e})>t(\mathbf{c})$ whenever $\lVert\Phi(\mathbf{c}^{\prime},\mathbf{e})-\Phi(\mathbf{c}^{\prime},\mathbf{e}^{\prime})\rVert$ is first-order stochastic dominant over $\lVert\Phi(\mathbf{c},\mathbf{e}^{\prime})-\Phi(\mathbf{c}^{\prime},\mathbf{e}^{\prime})\rVert$ with $\mathbf{c}^{\prime}\sim\mathcal{C},\mathbf{e}^{\prime}\sim\mathcal{E}$ uniformly.
@@ 4.2 De-biasing by Time-steps
1 2 4 8 16 1 2 4 8 16
CLIP Zero-Shot [26] 24.9 12.5
CoOp [48] 22.8 -2.1 28.4 +3.5 32.4 +7.5 37.7 +12.8 40.5 +15.6 10.8 -1.7 19.3 +6.8 19.6 +7.1 24.2 +11.7 25.8 +13.3
Co-CoOp [47] 30.1 +5.2 31.6 +6.7 33.6 +8.7 37.3 +12.4 38.2 +13.3 11.4 -1.1 13.9 +1.4 16.6 +4.1 21.8 +9.3 22.8 +10.3
MaPLe* [15] 30.1 +5.2 33.0 +8.1 33.8 +8.9 39.4 +14.5 40.7 +15.8 11.8 -0.7 14.5 +2.0 18.0 +5.5 22.6 +10.1 26.8 +14.3
OpenCLIP Zero-Shot [26] 42.3 16.9
Linear-probe [26] 18.4 -23.9 32.5 -9.8 44.1 +1.8 55.0 +12.7 59.8 +17.5 12.4 -4.5 14.5 -2.4 17.7 +0.8 20.1 +3.2 21.6 +4.7
Tip-Adapter [44] 47.7 +5.4 51.6 +9.3 54.7 +12.4 58.4 +16.0 62.2 +19.9 22.6 +5.7 25.3 +8.4 23.6 +6.7 31.1 +14.2 33.9 +17.0
Tip-Adapter-F [44] 48.4 +6.1 53.9 +11.6 56.9 +14.7 62.0 +19.7 67.4 +25.1 21.4 +4.5 22.3 +5.4 26.8 +9.9 34.4 +17.5 40.3 +23.4
CaFo* [45] 51.7 +9.4 54.6 +12.3 58.5 +16.2 63.2 +20.9 66.7 +24.4 - - - - -
DM Zero-Shot [17] 24.3 11.7
@@ 5.2 Main Results
1 2 4 8 16 1 2 4 8 16
CLIP CoOp [48] 8.2 10.4 17.2 29.6 33.7 11.4 13.9 18.1 30.3 34.1
Co-CoOp [47] 9.7 12.1 20.1 32.7 44.5 30.1 31.6 33.6 37.3 38.2
MaPLe [15] 13.5 32.9 40.7 55.7 63.0 35.2 40.7 44.4 57.5 68.1
OC Linear-probe [26] 11.9 13.2 37.7 53.5 60.2 16.7 29.5 51.1 61.1 69.8
Tip-Adapter [44] 28.5 36.9 46.5 59.3 66.7 36.4 47.5 59.8 71.2 80.1
Tip-Adapter-F [44] 29.3 32.0 54.4 74.2 82.6 37.4 48.6 62.2 79.1 85.4
@@
1 2 4 8 16 1 2 4 8 16
Subset last 48.5 55.8 64.2 73.1 78.1 36.9 53.6 73.1 84.2 91.3
last + $w_{1}$ 48.2 55.7 65.0 74.2 78.5 28.7 47.7 69.7 84.5 91.1
last + $w_{1}$+$w_{0}$ 46.5 55.4 65.1 74.0 79.9 25.8 43.2 65.7 83.4 91.6
Weight ELB weight 44.7 50.2 59.9 68.7 71.2 33.2 50.2 68.5 78.3 81.8
PDAE 49.2 56.7 64.1 72.2 78.1 36.2 53.9 73.4 81.2 90.8
Ours 48.5 55.8 64.2 74.2 79.9 36.9 53.6 73.1 84.5 91.6
@@
Overall Results. As shown in Table 1 and 2, our TiF learner achieves state-of-the-art performance across fine-grained and reID datasets, e.g., significantly improving existing methods by 13.7% on FGVCAircraft, 21.6% on DukeMTMC-reID and 16% on VeRi-776. On ISIC2019, we still achieve up to 7% absolute gain over the best-performing baseline, despite the challenges posed by diverse appearances within each class, as well as the stringent macro-F1 evaluation, which is sensitive to under-performing classes. Note that prompt tuning methods generally have non-ideal performances. This holds especially on customized reID classes with no class name to provide semantic prior. This validates the difficulties to describe the specification of fine-grained classes by prompt alone.
@@ 6 Conclusions
7 Acknowledgement
This research is supported by the National Research Foundation, Singapore under its AI Singapore Programme (AISG Award No: AISG2-RP-2021-022), MOE AcRF Tier 2 (MOE2019-T2-2-062), Wallenberg-NTU Presidential Postdoctoral Fellowship, the A*STAR under its AME YIRG Grant (Project No.A20E6c0101) and the Lee Kong Chian (LKC) Fellowship fund awarded by Singapore Management University. Pan Zhou was supported by the Singapore Ministry of Education (MOE) Academic Research Fund (AcRF) Tier 1 grant.